# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ 4.2 Complexities for ($\mathcal{P}$) and discussions
Comparison with communication lower bounds. When $S=1$ (and so $\overline{r^{2}}=(\overline{r})^{2}$), $K=I-\Pi$ (and so $\|K\|=\sigma_{\min}^{+}(K)$), with $w^{x}(x)=\frac{1}{2}\|x\|^{2}$ and $w^{y}_{s}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$, assuming that $\|f_{v}^{\prime}\|\leq M_{f}$ (and take $M=2M_{f}$), $a_{1}^{2}\leq m\gamma^{2}$ for some $\gamma\leq M_{f}$, we have $N=O(\frac{\overline{r}\cdot\sqrt{m}\gamma}{\sqrt{\epsilon\mu}})$. Since communication is only needed when $\mathsf{Agent}(y_{1})$ updates, the total number of communication rounds is $N/r_{s}=O(\frac{\sqrt{m}\gamma}{\sqrt{\epsilon\mu}})$, which achieves the theoretical lower bound (Theorem 2 and discussion after in [ArjevaniShamir2015Communication]) on the communication round complexity for $\mu$-strongly convex, $\sqrt{m}\gamma$-similar functions, and so is optimal555[ArjevaniShamir2015Communication] constructs a pair of “chain like” functions $\{\gamma F_{1},\gamma F_{2}\}$ which are $\Theta(\gamma)$-similar and $\mu$-strongly convex. In addition, when $m/2$ agents are given $\gamma F_{1}$ and the rest are given $\gamma F_{2}$ (and so this set of $m$ functions is $\Theta(\sqrt{m}\gamma)$-similar), the number of rounds of communication needed is $\Omega(\gamma\sqrt{\frac{1}{\mu\epsilon/m}})$..
5 Numerical experiments: support vector machine
@@
We consider the Support Vector Machine (SVM) problem with hinge loss and additional regularization. More precisely, each primal $\mathsf{Agent}(x_{v})$ is given $m_{s}$ pairs $\{(b_{v}^{l},y_{v}^{l})\}_{l\in[m_{s}]}$ such that $b_{v}^{l}\in\mathbb{R}^{d}$ is a feature vector satisfying $\|b_{v}^{l}\|=1$, and $y_{v}^{l}\in\{\pm 1\}$ is the label. The goal of SVM is to find a weight vector $x\in\mathbb{R}^{d}$ such that the linear classifier $b\to\mathsf{sign}(\langle b,x\rangle)$ agrees with most pairs $(b_{v}^{l},y_{v}^{l})$ in the dataset. To achieve this, one common approach is to solve the following (regularized) hinge loss minimization problem (in a distributed fashion)[lan_communication-efficient_2020, Duchi2012DA]:
$\min_{x\in\mathcal{X}}\sum_{v\in V}f_{v}(x),\quad f_{v}(x)=\frac{1}{m_{s}}\sum_{l=1}^{m_{s}}[1-y_{v}^{l}\langle b_{v}^{l},x\rangle]_{+}+\frac{\mu}{2}\|x\|^{2},~{}v\in V,$ (23)
@@
where we take $d=50$, $\mathcal{X}=\{x\in\mathbb{R}^{d}|\|x\|\leq 5\}$, and $[c]_{+}=\max(0,c)$ for all $c\in\mathbb{R}$. When $\mu=0$, (23) is the classical hinge loss minimization problem, and when $\mu>0$, the local objectives are $\mu$-strongly convex.
Setup. We take $w^{x}(x)=\frac{1}{2}\|x\|^{2}$ and $w^{y}_{s}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$, and
@@
Since $\|b_{v}^{l}\|\leq 1$ and $\|x\|\leq 5$ (since $x\in\mathcal{X}$), we have $\|f_{v}^{\prime}\|\leq 1+5\mu$ and so we can take $M=2(1+5\mu)$.
Below, we look at the suboptimality of $F(\Pi\underline{X}^{k})$ where $\underline{X}^{k}:=\frac{\sum_{k^{\prime}=0}^{k}\theta_{k^{\prime}}\widehat{X}^{k^{\prime}}}{\sum_{k^{\prime}=0}^{k}\theta_{k^{\prime}}}$. In Section 5.1, we focus on how $F(\Pi\underline{X}^{k})$ depends on the iteration $k$ and the mean updating rates $\overline{r}$. In Section 5.2, we focus on how $F(\Pi\underline{X}^{k})$ depends on the iteration $k$ and function similarities $A$.
5.1 Suboptimality and $k,\overline{r}$
@@
Data simulation. We first sample $x^{*}\in\mathbb{R}^{d}$ on the unit sphere uniformly at random as the “true solution”. For each agent $\mathsf{Agent}(x_{v})$, we generate the feature vectors $b_{v}^{l}$ i.i.d. uniformly on the unit sphere and take $y_{v}^{l}=\mathsf{sign}(\langle b_{v}^{l},x^{*}\rangle)$ w.p. $1-p$, and $y_{v}^{l}=-\mathsf{sign}(\langle b_{v}^{l},x^{*}\rangle)$ otherwise, independent of $\{b_{v}^{l}\}_{v\in V,l\in[m_{s}]}$. We take $m_{s}=50$ and $p=0.05$.
@@
Communication setup. We consider the hierarchical setup, with $3$ layers of dual agents and $1$ layer of primal agents, where each non-leaf node has $5$ child nodes. Thus, there are $m=125$ primal agents and $S=31$ dual agents, and $|\mathsf{Chi}(s)|=5$ for each non-leaf node. We assume that the dual agents at layer $i$ are updated with rate $r_{i}$ for $i=1,2,3$, and test with various $(r_{1},r_{2},r_{3})$
[…]
@@
Results: $F(\Pi\underline{X}^{k})$ and $k$. In Figures 2(a) and 3(a), we present $F(\Pi\underline{X}^{k})$ as a function of $k$ for MT-GS and AMT-GS respectively. Different lines correspond to different $(r_{1},r_{2},r_{3},\overline{r})$, with the line colors indicating $\overline{r}$. As can be seen, for both MT-GS and AMT-GS, all settings of $(r_{1},r_{2},r_{3},\overline{r})$ converge or show trend of convergence, and the convergence is faster for smaller $\overline{r}$. In addition, comparing Figures 2(a) and 3(a), we see that strong convexity (with AMT-GS) indeed accelerates the convergence.
Results: $F(\Pi\underline{X}^{k})$ and $\overline{r}$. In Figure 2(b), we present $F(\Pi\underline{X}^{k})$ as a function of $\overline{r}$, taken at $k=400,800,\ldots,2800$ for MT-GS. In Figure 3(b), we present $F(\Pi\underline{X}^{k})$ as a function of $\overline{r}^{2}$, taken at $k=60,120,\ldots,480$ for AMT-GS. In both figures, the line colors indicate $k$. As can be seen, as $k$ increases, the suboptimaltiy of $F(\Pi\underline{X}^{k})$ is approximately linear in $\overline{r}$ for MT-GS and is approximately linear in $\overline{r}^{2}$ for AMT-GS, agreeing with our Theorem 3.1 and 4.1.
(a) $F(\Pi\underline{X}^{k})$ as a function of $k$; legends represent $(r_{1},r_{2},r_{3},\overline{r})$ and line colors represent $\overline{r}$.
(b) $F(\Pi\underline{X}^{k})$ as a function of $\overline{r}$; legends and line colors represent $k$.
@@
Figure 2: MT-GS ($\mu=0$), dependence of $F(\Pi\underline{X}^{k})$ on the iteration number $k$ and the mean updating rate $\overline{r}$.
(a) $F(\Pi\underline{X}^{k})$ as a function of $k$; legends represent $(r_{1},r_{2},r_{3},\overline{r})$ and line colors represent $\overline{r}$.
(b) $F(\Pi\underline{X}^{k})$ as a function of $\overline{r}^{2}$; legends and line colors represent $k$.
Figure 3: AMT-GS ($\mu=0.01$), dependence of $F(\Pi\underline{X}^{k})$ on the iteration number $k$ and the mean updating rate $\overline{r}$.
5.2 Suboptimality and $k,A$
@@
In this experiment, there is only one dual agent (thus $A=\Theta(a_{1})$) who updates at rate $r_{1}=1$. We focus on (normalized) $F(\Pi\underline{X}^{k})$ as a function of $k$ and function similarities $a_{1}$.
@@
Dataset induced function similarities. In this experiment, similarities between $\{f^{\prime}_{v}\}_{v\in V}$ are inherited from similarities in the local datasets. More precisely, we generate a global dataset $\{(b_{global}^{l},y_{global}^{l})\}_{l\in[m_{global}]}$ consisting of $m_{global}$ pairs of data. In addition, for each agent, we generate $m_{local}$ pairs of private data $\{(b_{local,v}^{l},y_{local,v}^{l})\}_{l\in[m_{local}]}$. Thus, $\mathsf{Agent}(x_{v})$ has access to $\{(b_{global}^{l},y_{global}^{l})\}_{l\in[m_{global}]}\cup\{(b_{local,v}^{l},y_{local,v}^{l})\}_{l\in[m_{local}]}$, a total of $m_{s}=m_{global}+m_{local}$ pairs of data.
Data simulation. We simulate the $x^{*}$ and all pairs of data the same as in Section 5.1, with $m_{s}=50$, $m_{local}=\lfloor\gamma m_{s}\rfloor$ and $m_{global}=m_{s}-m_{local}$, for $\gamma=0.1,0.2,\ldots,0.9$. We assume that there are $m=500$ primal agents. Due to the global dataset, we have $\|f^{\prime}_{v}-f^{\prime}_{v^{\prime}}\|\leq\frac{m_{local}}{m_{s}}$ for any $v,v^{\prime}\in V$, and so we can take $a_{1}=2\gamma\sqrt{m}$.
@@
  2. 2.
    Type-1 setup. In addition to the above $\gamma$-aware setup, we also test our algorithms for $R_{s}=100R_{s}^{prj}$, $a_{1}=2\sqrt{m}$, and $x_{v}^{init}=\mathbf{0}$, which we denote as type-1 setup. Compared to type-0, type-1 has larger dual domain size and ignores the function similarities, making it a closer approximation to the DCS algorithms in [lan_communication-efficient_2020].
For both types of setups, we test MT-GS for $N+1=500$ and $\mu=0$ and AMT-GS for $N+1=200$ and $\mu=0.01$. Since the datasets are different for different $\gamma$, below, for each $\gamma$, we normalize $F(\Pi\underline{X}^{k})$ such that $F(\mathbf{0})=m=500$ is normalized to $1$, and the minimum (over $k$ and two types) of $F(\Pi\underline{X}^{k})$ is normalized to 00.
Results: normalized $F(\Pi\underline{X}^{k})$ and $k$. In Figures 4(a) and 5(a), we present the normalized $F(\Pi\underline{X}^{k})$ as a function of $k$ for MT-GS and AMT-GS respectively. Different lines correspond to different types of setup and different $\gamma$, with the line colors indicating $\gamma$. As can be seen, our MT-GS and AMT-GS converge in all the tested settings, and strong convexity (with AMT-GS) accelerates the convergence.
Moreover, from Figure 4(a), solid curves are always below dotted curves of the same $\gamma$ (color), indicating that type-0 setups converge faster than type-1 setups. In addition, solid curves corresponding to smaller $\gamma$ (functions are more similar) are dominated by curves corresponding to larger ones, while dotted curves converge at roughly the same rates. These suggest that type-0 setups, with the $\gamma$-aware domain sizes and parameters, indeed take advantage of the function similarities to speed up the convergence.
Results: normalized $F(\Pi\underline{X}^{k})$ and $\gamma$. In Figure 4(b), we present normalized $F(\Pi\underline{X}^{k})$ as a function of $\overline{\gamma}$, taken at $k=100,200,300,400$ for MT-GS. The line colors indicate $k$. As can be seen, the normalized $F(\Pi\underline{X}^{k})$ is approximately linear in $\gamma$ for MT-GS, which agrees with our Theorem 3.1.
In Figure 5(b), we present normalized $F(\Pi\underline{X}^{k})$ as a function of $\gamma^{2}$, taken at $k=10,15,20,25,30,40,80,120,160$ for AMT-GS. The line colors indicate $k$. It appears that the normalized $F(\Pi\underline{X}^{k})$ is increasing in $\gamma^{2}$ for most $\gamma$ (the dip when $\gamma^{2}=0.36$ could be due to the randomness in the simulated dataset). This confirms that our AMT-GS can take advantage of function similarities. However, the normalized $F(\Pi\underline{X}^{k})$ does not appear to be linear in $\gamma^{2}$: for small $\gamma^{2}$, it does not converge to 00, which could be because setting $T=N+1$ (smaller than suggested) introduces additional suboptimality; in addition, it is possible that the normalization process introduces extra $\gamma$-dependent factors.
(a) Normalized $F(\Pi\underline{X}^{k})$ as a function of $k$; legends represent (type of $(\text{type of setup},\gamma)$ and line colors represent $\gamma$.
(b) Normalized $F(\Pi\underline{X}^{k})$ as a function of $\gamma$; legends represent (type of $(\text{type of setup},k)$ and line colors represent $k$.
Figure 4: MT-GS ($\mu=0$), dependence of normalized $F(\Pi\underline{X}^{k})$ on the iteration number $k$ and function similarities $\gamma$.
(a) Normalized $F(\Pi\underline{X}^{k})$ as a function of $k$; legends represent (type of $(\text{type of setup},\gamma)$ and line colors represent $\gamma$.
(b) Normalized $F(\Pi\underline{X}^{k})$ as a function of $\gamma^{2}$; legends represent (type of $(\text{type of setup},k)$ and line colors represent $k$.
Figure 5: AMT-GS ($\mu=0.01$), dependence of normalized $F(\Pi\underline{X}^{k})$ on the iteration number $k$ and function similarities $\gamma$.
6 Conclusion and future directions
In this work, we propose the (accelerated) multi-timescale gradient sliding algorithm for distributed optimization with general convex, non-smooth objectives. Our MT-GS and AMT-GS are deterministic, can take advantage of potential similarities between local objectives, and allow user-picked updating frequencies for different dual blocks. To make the algorithms more practical, one direction of future work is to develop (if possible) adaptive algorithms which, without prior knowledge of function similarities $\{a_{s}\}_{s\in[S]}$, still maintain communication round complexities which have optimal dependency on $\{a_{s}\}_{s\in[S]}$. Another direction of future work is to develop more space-efficient algorithms: currently each primal agents need to store $O(r_{\max})$ vectors in $\mathbb{R}^{d}$, which are used as mixtures of proximal centers and when calculating messages to dual agents. It is an interesting question whether one can achieve similar convergence guarantees with smaller memory requirement.

### v2

@@ 4.6 Convergence for strongly convex objectives
Comparison with communication lower bounds. When $S=1$ (and so $\overline{r^{2}}=(\overline{r})^{2}$), $K=I-\Pi$ (and so $\|K\|=\sigma_{\min}^{+}(K)$), with $w_{x}(x)=\frac{1}{2}\|x\|^{2}$ and $w_{y_{s}}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$, assuming that $\|f_{v}^{\prime}\|\leq M_{f}$, $a_{1}^{2}\leq m\gamma^{2}$ for some $\gamma\leq M_{f}$, we have $N=O(\frac{\overline{r}\cdot\sqrt{m}\gamma}{\sqrt{\epsilon\mu_{f}}})$. Since communication is only needed when $\mathsf{Agent}(y_{1})$ updates, the total number of communication rounds is $N/r_{s}=O(\frac{\sqrt{m}\gamma}{\sqrt{\epsilon\mu_{f}}})$, which achieves the theoretical lower bound (Theorem 2 and discussion after in [4]) on the communication round complexity for $\mu$-strongly convex, $\sqrt{m}\gamma$-similar functions, and so is optimal666[4] constructs a pair of “chain like” functions $\{\gamma F_{1},\gamma F_{2}\}$ which are $\Theta(\gamma)$-similar and $\mu$-strongly convex. In addition, when $m/2$ agents are given $\gamma F_{1}$ and the rest are given $\gamma F_{2}$ (and so this set of $m$ functions is $\Theta(\sqrt{m}\gamma)$-similar), the number of rounds of communication needed is $\Omega(\gamma\sqrt{\frac{1}{\mu_{f}\epsilon/m}})$..
5 Numerical experiments
Below, we present numerical experiments applying MT-PDHG to linear programming problems (Section 5.1), and (A)MT-PDHG to distributed Support Vector Machine problems (Section 5.2). All experiments are implemented using Python and run on MacBook Air with the M3 chip and 8 cores.
5.1 Experiment: linear programming
In this set of experiments, we apply our MT-PDHG and the vanilla PDHG to simulated linear programming problems of the form
$\min_{X\in\mathbb{R}^{n}}c^{T}X,\quad s.t.~AX=b,~X\geq\mathbf{0},$
where $A\in\mathbb{R}^{m\times n}$ and the rows are divided evenly into $S=6$ blocks, with the associated dual blocks updated in parallel:
$\min_{X\in\mathbb{R}^{n}}\max_{Y=(y_{s})_{s\in[S]}\in\mathbb{R}^{m}}c^{T}X-\sum_{s=1}^{S}(y_{s}^{T}A_{s}X-y_{s}^{T}b),\quad s.t.~X\geq\mathbf{0}.$
Problem simulation. We simulate $c_{i}\sim\mathcal{N}(0,1)$ for $i\in[n]$, $A_{i,j}\sim\text{Uniform}([0,1))$ for $i\in[m],j\in[n]$, all independently. To ensure the problem is feasible, we simulate $X^{\prime}_{i}\sim\text{Uniform}([0,1))$ for $i\in[n]$ independent of $c,A$, and take $b=AX^{\prime}$.
Choice of rates $r_{s}$. We consider 4 combinations of the updating rates for the dual blocks: 1. $r_{s}=1$ for all $s$; 2. $r_{1}=r_{2}=r_{3}=1$ and $r_{4}=r_{5}=r_{6}=10$; 3. $r_{s}=10$ for all $s$; 4. $r_{s}=50$ for all $s$. These are denoted in different colors in Figures 3 and 4.
Algorithm setups. We use $w_{X}(X)=\frac{1}{2}\|X\|^{2}$ and $w_{y_{s}}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$. We benchmark our MT-PDHG against the vanilla PDHG, where at each global iteration $k$, $X$ is updated as the minimizer to $\langle c-A^{T}\overline{Y}^{k},X\rangle+\frac{\eta}{2}\|X-X^{k-1}\|^{2}$, with $\eta=\|A\|$, and at $k=i_{s}r_{s}$, the dual block $y_{s}$ is updated as the minimizer to $\langle K_{s}\widetilde{X}_{s}^{i_{s}}-b,y_{s}\rangle+\frac{\tau_{s}}{2}\|y_{s}-y_{s}^{i_{s}-1}\|^{2}$, where $\widetilde{X}_{s}^{i_{s}}=2X^{k-1}-X^{k-2}$ and $\tau_{s}=2S\|A_{s}\|^{2}/\eta$. Thus, there is no mixture of Bregman divergence and multi-timescale extrapolation. For our MT-PDHG, we use $\rho_{s}=1/S$ and the same $\eta,\tau_{s}$ as above. We initialize $X=\mathbf{0}$ and $Y=\mathbf{0}$.
Results. We present the KKT residual $(\|AX-b\|^{2}+\|[A^{T}Y-c]_{+}\|^{2}+[c^{T}X-b^{T}Y]_{+})^{1/2}$ for MT-PDHG and vanilla PDHG, under different combinations of $r_{s}$.
In Figure 3, we present the residual as a function of global iteration. As can be seen, our MT-PDHG is stable under various combinations of updating rates. Interestingly, the vanilla PDHG, even without mixture of Bregman divergence and multi-timescale extrapolation, still converges when the rates are small (rate combinations 1, 2, and 3). However, as the yellow lines suggest, when the rates ($r_{s}=50$) are large, our explicit control through the Bregman divergence and the extrapolation helps stabilize the performance. Moreover, comparing the green and blue curves, which have the same $\max_{s}r_{s}=10$ but different average $r_{s}$, which is $5.5$ for green but $10$ for blue, we see that smaller $\overline{r}$ indeed corresponds to faster convergence rate (as a function of iteration), demonstrating that our MT-PDHG are robust to extreme values in $r_{s}$.
In Figure 4, we rescale the $x$-axis of Figure 3 by the total runtime of each configuration. Comparing MT-PDHG with vanilla PDHG, we observe an additional overhead arising from the computation of the Bregman mixture and multi-timescale extrapolation, which involves evaluating $\sum_{s}\rho_{s}X^{k-r_{s}}$ and $\sum_{i=1}^{r_{s}}X^{k-i}$. However, as $m$ and $n$ grow, this overhead becomes negligible since the runtime is increasingly dominated by matrix–vector multiplications. We also observe that, approximately, the wall-clock time required to reach a target accuracy scales with $\max_{s}r_{s}$ (instead of $\overline{r}$ for the iteration). We note, however, that this behavior may depend on implementation details and the underlying computing hardware.
(a) $(m,n)=(500,1000)$
(b) $(m,n)=(1000,3000)$
(c) $(m,n)=(3000,5000)$
Figure 3: KKT residual as a function of global iteration for different $(m,n)$, under 4 different combinations of the updating rates for the dual blocks: 1. $r_{s}=1$ for all $s$; 2. $r_{1}=r_{2}=r_{3}=1$ and $r_{4}=r_{5}=r_{6}=10$; 3. $r_{s}=10$ for all $s$; 4. $r_{s}=50$ for all $s$.
(a) $(m,n)=(500,1000)$
(b) $(m,n)=(1000,3000)$
(c) $(m,n)=(3000,5000)$
Figure 4: KKT residual as a function of running time in seconds for different $(m,n)$, under 4 different combinations of the updating rates for the dual blocks: 1. $r_{s}=1$ for all $s$; 2. $r_{1}=r_{2}=r_{3}=1$ and $r_{4}=r_{5}=r_{6}=10$; 3. $r_{s}=10$ for all $s$; 4. $r_{s}=50$ for all $s$.
5.2 Experiment: distributed Support Vector Machine
@@
We consider the Support Vector Machine (SVM) problem with hinge loss and additional regularization. More precisely, each primal $\mathsf{Agent}(x_{v})$ is given $m_{s}$ pairs $\{(b_{v}^{l},y_{v}^{l})\}_{l\in[m_{s}]}$ such that $b_{v}^{l}\in\mathbb{R}^{\overline{d}}$ is a feature vector satisfying $\|b_{v}^{l}\|=1$, and $y_{v}^{l}\in\{\pm 1\}$ is the label. The goal of SVM is to find a weight vector $x\in\mathbb{R}^{\overline{d}}$ such that the linear classifier $b\to\mathsf{sign}(\langle b,x\rangle)$ agrees with most pairs $(b_{v}^{l},y_{v}^{l})$ in the dataset. To achieve this, one common approach is to solve the following (regularized) hinge loss minimization problem (in a distributed fashion)[27, 19]:
$\min_{x\in\overline{\mathcal{X}}}\sum_{v\in V}f_{v}(x),\quad f_{v}(x)=\frac{1}{m_{s}}\sum_{l=1}^{m_{s}}[1-y_{v}^{l}\langle b_{v}^{l},x\rangle]_{+}+\frac{\mu}{2}\|x\|^{2},~v\in V,$ (31)
@@
In this experiment, we use the w8a dataset in LIBSVM [16], which consists of $49749$ samples, and the feature dimension is $\overline{d}=300$. We first normalize the features $\|b_{v}^{l}\|=1$ for all data, and take $\overline{\mathcal{X}}=\{x\in\mathbb{R}^{\overline{d}}|\|x\|\leq 5\}$. When $\mu=0$, (31) is the classical hinge loss minimization problem, and when $\mu>0$, the local objectives are $\mu$-strongly convex.
Setup. We take $w_{x}(x)=\frac{1}{2}\|x\|^{2}$ and $w_{y_{s}}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$, and
@@
Since $\|b_{v}^{l}\|\leq 1$ and $\|x\|\leq 5$ (since $x\in\overline{\mathcal{X}}$), we have $\|f_{v}^{\prime}\|\leq 1+5\mu$ and so we can take $M=2(1+5\mu)$.
Below, we look at the suboptimality of $F(\Pi\underline{X}^{k})$ where $\underline{X}^{k}:=\frac{\sum_{k^{\prime}=0}^{k}\theta_{k^{\prime}}\widehat{X}^{k^{\prime}}}{\sum_{k^{\prime}=0}^{k}\theta_{k^{\prime}}}$.
5.2.1 Suboptimality and $k,\overline{r}$
@@
Communication setup. We consider the hierarchical setup, with $3$ layers of dual agents and $1$ layer of primal agents, where each non-leaf node has $5$ child nodes. Thus, there are $m=125$ primal agents and $S=31$ dual agents, and $|\mathsf{Chi}(s)|=5$ for each non-leaf node. We assume that the dual agents at layer $i$ are updated with rate $r_{i}$ for $i=1,2,3$, and test with various $(r_{1},r_{2},r_{3})$. We divide the data set evenly among the primal agents.
Algorithm setup. Notice that the $\mu x$ part in $f^{\prime}_{v}$ is common to all $v$ so we can take $a_{s}=2\sqrt{|\mathsf{Des}(s)|}$ for all $s$. We use $R_{s}^{prj}$ as defined in (21) with $\xi=1$, and set $x_{v}^{init}=\mathbf{0}$ and $y_{s}^{init}=\mathbf{0}$. We test MT-PDHG for $N+1=3000$ and $\mu=0$. All parameters are set according to Theorem 3.1. We test AMT-PDHG for $N+1=500$ and $\mu=0.01$. We set $T=N+1$ for simplicity, and set all other parameters according to Theorem 3.2.
Results. In Figures 5, we present $F(\Pi\underline{X}^{k})-F^{*}$ as a function of $k$ for MT-PDHG and AMT-PDHG respectively, where $F^{*}$ is the minimum value among all iterations of all algorithm configurations plus $0.001$ (as the $y$-axis is in the $\log$ scale).
Different lines correspond to different $(r_{1},r_{2},r_{3},\overline{r})$, with the line colors indicating $\overline{r}$, and the line with starred markers can serve as the benchmark: when all $r_{s}=1$, our (A)MT-PDHG has the same updating rules as the classical PDHG.
@@
From the figures, we see that under all settings of $(r_{1},r_{2},r_{3},\overline{r})$, our (A)MT-PDHG converge or show trend of convergence, and the convergence is faster for smaller $\overline{r}$. In addition, comparing the two figures in Figure 5, we see that strong convexity (with AMT-PDHG) indeed accelerates the convergence.
@@
Figure 5: Dependence of $F(\Pi\underline{X}^{k})$ on the iteration number $k$ and the mean updating rate $\overline{r}$ for MT-PDHG ($\mu=0$) (left) and AMT-PDHG ($\mu=0.01$) with communication sliding. Legends represent $(r_{1},r_{2},r_{3},\overline{r})$ and line colors represent $\overline{r}$.
5.2.2 Suboptimality and costs of communication
To further demonstrate the potential benefit of using different updating rates, we rescale the lines in the left plots of Figure 5 based on Amortized Costs ($AC$): since each of the $5^{i-1}$ dual agents at layer $i$ is updated every $r_{i}$ global iteration and each iteration costs $c_{i}$, the average cost per global iteration becomes $AC:=c_{1}/r_{1}+5c_{2}/r_{2}+25c_{3}/r_{3}$. In Figures 6 and 7, we rescale the $x$-axis in Figure 5 using $AC$ for each combinations of $r_{s}$, and the resulting plots reflect the objective values as a function of the costs. In addition, we use different markers to represent the ratio $r_{1}:r_{2}:r_{3}$: circle for $1:1:1$, triangle for $3:2:1$, and star for $1:2:3$.
From Figure 6, we see that the convergence rate for fixed $(c_{1},c_{2},c_{3})$ depends on the ratios of $(r_{1},r_{2},r_{3})$, as the lines with the same ratio (marker type) coincide. In addition, when the costs vary significantly, there is significant benefit of choosing the rates adaptive to the costs. For instance, in the middle two figures, with large $c_{1}$ or $c_{2}$, MT-PDHG is more cost-efficient when $r_{1},r_{2}$ are large (as then dual agents in top two layers of the tree are updated less frequently). Indeed, lines with star markers show a faster rate of convergence.
For AMT-PDHG, as suggested by our convergence results in Corollary 4.4, the convergence rates have a more complicated dependency on $r_{s}$. From Figure 7, the relative level of $(r_{1},r_{2},r_{3})$ does not determine the convergence rates any more. However, the results still agree with the general intuition that with larger $c_{1}$ or $c_{2}$, it’s more cost-efficient to choose relatively large $r_{1}$ or $r_{2}$.
Figure 6: MT-PDHG ($\mu=0$), dependence of $F(\Pi\underline{X}^{k})$ on the costs. Legend represents $(r_{1},r_{2},r_{3},\overline{r},AC)$ where the amortized cost $AC:=c_{1}/r_{1}+5c_{2}/r_{2}+25c_{3}/r_{3}$. The marker indicates the ratio $r_{1}:r_{2}:r_{3}$: circle for $1:1:1$, triangle for $3:2:1$, and star for $1:2:3$.
Figure 7: AMT-PDHG ($\mu=0.01$), dependence of $F(\Pi\underline{X}^{k})$ on the costs. Legend represents $(r_{1},r_{2},r_{3},\overline{r},AC)$ where the amortized cost $AC:=c_{1}/r_{1}+5c_{2}/r_{2}+25c_{3}/r_{3}$. The marker indicates the ratio $r_{1}:r_{2}:r_{3}$: circle for $1:1:1$, triangle for $3:2:1$, and star for $1:2:3$.
5.2.3 Suboptimality and $k,A$
@@
In this experiment, we use $1\%$ of the w8a dataset. We assume that there is only one dual agent (thus $A=\Theta(a_{1})$) who updates at rate $r_{1}=1$, and there are $m=10$ primal agents. We focus on (normalized) $F(\Pi\underline{X}^{k})$ as a function of $k$ and function similarities $a_{1}$.
@@
Dataset induced function similarities. In this experiment, similarities between $\{f^{\prime}_{v}\}_{v\in V}$ are inherited from similarities in the local datasets. More precisely, there is a global dataset $\{(b_{global}^{l},y_{global}^{l})\}_{l\in[m_{global}]}$ consisting of $m_{global}$ pairs of data. In addition, each agent has $m_{local}$ pairs of private data $\{(b_{local,v}^{l},y_{local,v}^{l})\}_{l\in[m_{local}]}$. Thus, $\mathsf{Agent}(x_{v})$ has access to $\{(b_{global}^{l},y_{global}^{l})\}_{l\in[m_{global}]}\cup\{(b_{local,v}^{l},y_{local,v}^{l})\}_{l\in[m_{local}]}$, a total of $m_{s}=m_{global}+m_{local}$ pairs of data.
More concretely, the global dataset consists of $\frac{1-\gamma}{1+(m-1)\gamma}$ fraction of the data, and the rest of the data is divided evenly among $m$ primal agents as local data. Thus $m_{local}\approx\gamma m_{s}$. We test $\gamma=0.1,0.2,\ldots,0.8$. Due to the global dataset, we have $\|f^{\prime}_{v}-f^{\prime}_{v^{\prime}}\|\leq\frac{m_{local}}{m_{s}}$ for any $v,v^{\prime}\in V$, and so we take $a_{1}=2\gamma\sqrt{m}$.
@@
    Type-0 setup. We use $R_{s}=R_{s}^{prj}$ as defined in (21) with $\xi=1$, and $y_{s}^{init}=\mathbf{0}$. We set the parameter $T=N+1$ and all other parameters are set according to Theorems 3.1 and 3.2. For the initialization, for MT-PDHG, we use $x_{v}^{init}=\mathbf{0}$ and for AMT-PDHG, we use $\underline{x}_{v}^{0}=\mathbf{0}$ and construct $x_{v}^{init}$ according to Corollary 4.3.
  2. 2.
    Type-1 setup. In addition to the above $\gamma$-aware setup, we also test our algorithms for $R_{s}=10000R_{s}^{prj}$, $a_{1}=2\sqrt{m}$, and $x_{v}^{init}=\mathbf{0}$, which we denote as type-1 setup. Compared to type-0, type-1 has larger dual domain size and ignores the function similarities. Thus, it can serve as an approximation to the DCS algorithms in [27].
For both types of setups, we test MT-PDHG for $N+1=500$ and $\mu=0$ and AMT-PDHG for $N+1=200$ and $\mu=0.01$. Since the datasets are different for different $\gamma$, below, for each $\gamma$, we normalize $F(\Pi\underline{X}^{k})$ such that $F(\mathbf{0})=m=500$ is normalized to $1$, and the minimum (over $k$ and two types) of $F(\Pi\underline{X}^{k})$ is normalized to 0.
Results. In Figure 8, we present the normalized $F(\Pi\underline{X}^{k})$ as a function of $k$ for MT-PDHG and AMT-PDHG respectively. Different lines correspond to different types of setup and different $\gamma$, with the line colors indicating $\gamma$. As can be seen, our MT-PDHG and AMT-PDHG converge in all the tested settings, and strong convexity (with AMT-PDHG) accelerates the convergence.
Moreover, for MT-PDHG, from the left figure in Figure 8 solid curves – representing $\gamma$-aware setup – are converging faster than the dotted curves, In addition, as $\gamma$ decreases – global dataset takes a larger fraction – the algorithm converges faster. These demonstrate that our proposed function similarity dependent penalties indeed take advantage of the function similarities to speed up the convergence.
For the AMT-PDHG, from the right figure in Figure 8, one can see that similarity helps speed up the convergence for both types of setups. We leave it to future works to investigate if the conditions on the penalty levels are necessary for function-similarity dependent convergence rates when the local objectives are strongly convex.
@@
6 Conclusion and future direction
In this work, we propose the (accelerated) multi-timescale PDHG algorithms for saddle point problems with block-decomposable duals. Our (A)MT-PDHG allows arbitrary updating rates for dual blocks while remaining fully deterministic and robust to extreme delays in dual updates. We further apply (A)MT-PDHG to distributed optimization and demonstrate how the flexibility in choosing the updating rates could help improve the overall algorithm efficiencies in heterogeneous environments.
To make the algorithms more practical, one direction of future work is to develop more space-efficient algorithms: currently each primal agents need to store $O(r_{\max})$ vectors in $\mathbb{R}^{d}$, which are used as mixtures of proximal centers and when calculating messages to dual agents. It is an interesting question whether one can achieve similar convergence guarantees with smaller memory requirement. Another promising direction is to extend the multi-timescale update mechanism to a broader class of algorithms, including those for saddle point and more general optimization problems.